# Participation bias in the estimation of heritability and genetic correlation

### v2

@@ Abstract
It is increasingly recognized that participation bias can pose problems for genetic studies. Recently, to overcome the challenge that genetic information of non-participants is unavailable, it is shown that by comparing the IBD (identity by descent) shared and not-shared segments among the participants, one can estimate the genetic component underlying participation. That, however, does not directly address how to adjust estimates of heritability and genetic correlation for phenotypes correlated with participation. Here, for phenotypes whose mean differences between population and sample are known, we demonstrate a way to do so by adopting a statistical framework that separates out the genetic and non-genetic correlations between participation and these phenotypes. Crucially, our method avoids making the assumption that the effect of the genetic component underlying participation is manifested entirely through these other phenotypes. Applying the method to 12 UK Biobank phenotypes, we found 8 have significant genetic correlations with participation, including body mass index, educational attainment, and smoking status. For most of these phenotypes, without adjustments, estimates of heritability and the absolute value of genetic correlation would have underestimation biases.
@@ 1 Main
For sample surveys in general, a common approach to adjusting for ascertainment bias is to construct a propensity score based on variables, e.g., sex and educational attainment, whose distributional differences between sample and target population are known, or can be well approximated 8. By assuming that the systematic component of participation probability is fully captured by this score, analyses can be adjusted by applying inverse propensity weighting (IPW) to the samples. For genetic studies, it has been shown that participation can be associated with many phenotypes, such as educational attainment, alcohol use, mental and physical health 9; 10; 11; 12; 13; 14; 15. The UKBB participants are reported to be less likely to be obese, to smoke, to drink alcohol daily, and to have fewer self-reported health conditions compared with the general UK population.13. A recent study applied IPW to the UKBB data by creating a propensity score based on 14 phenotypes including age, BMI, weight, education, etc. 15 This propensity score, however, does not include genotypes. Thus, for analyses that involve genotypes, this adjustment would be sufficient only under the assumption that the systematic genotypic differences between sample and population are fully captured by this propensity score, which is equivalent to saying that the effect of the genetic component underlying participation is manifested entirely through this score, which can be considered as a composite phenotype. Previous results16 and results presented below show that this assumption is unlikely to hold.
@@
While the effectiveness of IPW or any other adjustment methods based only on phenotypes is doubtful, there is no obvious alternative unless genotype difference between sample and population can be independently estimated without relying on the phenotypes, which is difficult given that genotypes of non-participants are unavailable. A recent publication showed that the sample-population allelic frequency differences can be estimated by comparing the IBD (identity by descent) shared and not-shared segments among the participants. Here we utilize information obtained from that method through introducing a statistical model that specifies a genetic and a non-genetic component for each of the participation variable and the other phenotypes. The model separates out the genetic and non-genetic correlations between participation and other correlated phenotypes, and allows us to obtain adjusted estimates of heritability and genetic correlation for these phenotypes that take PB into account.
@@
This model includes the IPW model as a special case. Theoretically, without adjustments, heritability and genetic correlation can be over-estimated or under-estimated depending on the relative magnitudes of the genetic and non-genetic correlations between participation and the phenotypes. Empirically, we applied the adjustment method to 12 phenotypes of the UKBB data, and found that, without adjustment, heritability and the absolute value of the genetic correlation estimates tend to be underestimated for most of them.
@@ 2.1 Participation model overview
In practice, participation is often a two-step process. In the first step, a group of individuals are invited to participate in the study, and in the second step the invited make the decision whether to participate. For simplicity, we consider a model where the bias is only in the second step, $i.e.$ the invited list is representative of the target population. Consequences of the violation of this assumption are discussed later. For the second step, as in Benonisdottir and Kong (2023)16, we adopt a liability-threshold model where the liability score of a person is denoted by $X$, which is assumed to have (approximately) a standard normal distribution in the population. An invited person participates if $X\geq\Phi^{-1}(1-\alpha)$, where $\Phi$ is the standard normal cumulative distribution function and $\alpha$ is the participation rate of those invited. A phenotype of interest is denoted by $Y$, standardized to have mean zero and variance 1. For an individual in the invited list, we assume it is a random draw following the additive model:
$X=\boldsymbol{G}\boldsymbol{\beta}+{\epsilon}_{x},$ (1)
$Y=\boldsymbol{G}\boldsymbol{\gamma}+{\epsilon}_{y},$
@@
where $\boldsymbol{G}$ is the row vector of standardized genotypes; $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ are the per-standardized genotype effect sizes; and ${\epsilon}_{x}$ and ${\epsilon}_{y}$ are parts of $X$ and $Y$ that are orthogonal (uncorrelated) to the genetic components $\boldsymbol{G}\boldsymbol{\beta}$ and $\boldsymbol{G}\boldsymbol{\gamma}$, which could be partly completely random and partly determined by non-genetic factors. We refer to ${\epsilon}_{x}$ and ${\epsilon}_{y}$ as the non-genetic factors even though in practice they might also include non-additive genetic effects and effects of genotypes not included in $\boldsymbol{G}$. We denote the narrow-sense heritability of $X$ and $Y$ in the population as $h^{2}_{x}$ and $h^{2}_{y}$, which measure the phenotypic variance explained by additive effects of genetic components. The genetic and non-genetic correlations between participation liability score $X$ and phenotype $Y$ in the population are denoted as $\rho_{g}=\mathrm{Corr}(\boldsymbol{G}\boldsymbol{\beta},\boldsymbol{G}\boldsymbol{\gamma})$ and $\rho_{e}=\mathrm{Corr}({\epsilon}_{x},{\epsilon}_{y})$. Correlation between $X$ and $Y$, denoted by $\rho$, is equal to $\rho=\rho_{g}\sqrt{h^{2}_{x}h^{2}_{y}}+\rho_{e}\sqrt{(1-h^{2}_{x})(1-h^{2}_{y})}$.
@@ 2.2 Heritability estimated with participation bias
Under model (1), the heritability of phenotype $Y$ can be calculated as $h^{2}_{y}=\mathrm{Corr}(Y,\boldsymbol{G}\boldsymbol{\gamma})^{2}=\boldsymbol{\gamma}^{T}\bm{R}\boldsymbol{\gamma}$, where $\bm{R}=\mathrm{Corr}(\boldsymbol{G})$ is the linkage disequilibrium (LD) matrix in the population. Let
$h^{2}_{y,PB}=\max_{\boldsymbol{\gamma}^{\prime}}\left[\mathrm{Corr}(Y,\boldsymbol{G}\boldsymbol{\gamma}^{\prime}|X\geq\Phi^{-1}(1-\alpha))^{2}\right],$ (2)
which is the “apparent” heritability that is estimated using the sample of participants without considering PB. Note that the $\boldsymbol{\gamma}^{\prime}$ maximising (2) is in general not proportional to $\boldsymbol{\gamma}$ (Supplementary Note 1.2), meaning that PB would create estimation bias for both the form and strength of the genetic component of $Y$. Many of the results presented below, in particular formulas (3), (5), and (6), are derived based on the assumption that genotypes and phenotypes jointly have a multivariate normal distribution (MVN)17 (see Supplementary Notes 1.1 and 1.2). While genotypes are obviously not normally distributed, the expressions derived below under the normal assumption, the approximate equalities derived, indicated by the $\simeq$ sign, are shown to be very good by simulations, when the number of variants that have an effect on the phenotype is large and each effect is small:
$h^{2}_{y,PB}\simeq\frac{1}{1-\xi(\alpha)\rho^{2}}\left[(h^{2}_{y}-2\xi(\alpha)\rho\rho_{G}+\xi(\alpha)^{2}\rho^{2}h^{2}_{x})+\frac{\xi(\alpha)}{1-\xi(\alpha)h^{2}_{x}}(\rho_{G}-\xi(\alpha)\rho h^{2}_{x})^{2}\right].$ (3)
Here $\rho_{G}=\rho_{g}\sqrt{h^{2}_{x}h^{2}_{y}}$ is the covariance of $\boldsymbol{G}\boldsymbol{\beta}$ and $\boldsymbol{G}\boldsymbol{\gamma}$; and $\xi(\alpha)$ is a monotonic decreasing function of the participation rate $\alpha$:
$\xi(\alpha)=\left[\frac{\phi\left[\Phi^{-1}(1-\alpha)\right]}{\alpha}\right]^{2}-\left[\Phi^{-1}(1-\alpha)\right]\cdot\left[\frac{\phi\left[\Phi^{-1}(1-\alpha)\right]}{\alpha}\right],$ (4)
where $\phi$ is the density function of the standard normal distribution. Notably, $\xi(\alpha)$ is a monotonically decreasing function in (0,1) with $\lim_{\alpha\to 0}\xi(\alpha)=1$, and $\lim_{\alpha\to 1}\xi(\alpha)=0$. By definition, the heritability is the proportion of phenotypic variance explained by the genetic component. In (3), the denominator can be regarded as the variance of $Y$ after selection, which is determined by the participation rate and the correlation between $X$ and $Y$. It is expected to shrink after selection. On the other hand, the numerator is more intricate as PB impacts both effect sizes of SNPs and the LD structure, and both contributes to the variance of the genetic component. Specifically, the term $-2\xi(\alpha)\rho\rho_{G}$ can be understood as the first-order contribution of the alterations in effect sizes to the genetic variance, and $\xi(\alpha)\rho^{2}_{G}/(1-\xi(\alpha)h^{2}_{x})$ comes from the alterations in LD structures. The remaining terms involve the joint effects of the alterations in effect sizes of $Y$ and the LD structure.
@@
A comparison of $h^{2}_{y}$ and $h^{2}_{y,PB}$ is shown in Figure 1(a). In general, with the other parameters fixed, the difference between $h^{2}_{y}$ and $h^{2}_{y,PB}$ increases as $\alpha$ decreases. It is worth noting that PB can lead to both upward and downward bias to the estimation of heritability in the sample of participants. For example, if the correlation between $X$ and $Y$ arises solely from the genetic components, conditioning on participation will reduce the variance of the genetic component of $Y$ in the selected sample. With the variance of the non-genetic component unchanged, the proportion of the genetic effects in the sample will shrink. Conversely, if the correlation between the phenotype and participation is induced by non-genetic factors only, the proportion of variance of $Y$ attributed to the genetic component in the sample will increase, which leads to an upward bias in the estimation of heritability.
@@
Figure 1: Theoretical results for heritability and genetic correlation in the sample under various genetic architectures. (a) The heritability of a phenotype of interest other than participation. (b) The genetic correlation between two phenotypes (other than participation). (c) The genetic correlation between the liability score of participation and the phenotype. The heritability of participation liability score and other phenotypes are fixed at 0.125 and 0.5. Dashed lines indicate population value and blue lines indicate what is being estimated based on the biased sample without adjustment.
@@ 2.3 Genetic correlation estimated with participation bias
Here we consider two phenotypes, $Y_{1}$ and $Y_{2}$, both following the additive model (1), and $\varphi_{g}$ denotes their genetic correlation in the population. The corresponding genetic covariance is $\varphi_{G}=\varphi_{g}\sqrt{h^{2}_{y_{1}}h^{2}_{y_{2}}}$, where $h^{2}_{y_{1}}$ and $h^{2}_{y_{2}}$ are the heritability of the two phenotypes. Let $\varphi_{g,PB}=\mathrm{Corr}(\boldsymbol{G}\boldsymbol{\gamma}^{\prime}_{1},\boldsymbol{G}\boldsymbol{\gamma}^{\prime}_{2}|X\geq\Phi^{-1}(1-\alpha))$, where $\boldsymbol{\gamma}^{\prime}_{1}$ and $\boldsymbol{\gamma}^{\prime}_{2}$ are per-standardized genetic effect sizes in the sample of participants, which maximize $\mathrm{Corr}(Y_{1},\boldsymbol{G}\boldsymbol{\gamma}^{\prime}_{1}|X\geq\Phi^{-1}(1-\alpha))^{2}$ and $\mathrm{Corr}(Y_{2},\boldsymbol{G}\boldsymbol{\gamma}^{\prime}_{2}|X\geq\Phi^{-1}(1-\alpha))^{2}$, respectively. In parallel with $h^{2}_{y,PB}$, the quantity $\varphi_{g,PB}$ represents the “apparent” correlation between the genetic components of $Y_{1}$ and $Y_{2}$ that is being estimated ignoring PB. We show in Supplementary Note 1.3 that,
$\varphi_{g,PB}\simeq\frac{\varphi_{G}-\xi(\alpha)(\rho_{1}\rho_{G_{2}}+\rho_{2}\rho_{G_{1}})+\xi(\alpha)^{2}\rho_{1}\rho_{2}h^{2}_{x}+\frac{\xi(\alpha)}{1-\xi(\alpha)h^{2}_{x}}(\rho_{G_{1}}-\xi(\alpha)\rho_{1}h^{2}_{x})(\rho_{G_{2}}-\xi(\alpha)\rho_{2}h^{2}_{x})}{\sqrt{(1-\xi(\alpha)\rho^{2}_{1})(1-\xi(\alpha)\rho^{2}_{2})}\cdot\sqrt{h^{2}_{y_{1},PB}h^{2}_{y_{2},PB}}}.$ (5)
Here $\rho_{G_{1}}$ and $\rho_{G_{2}}$ are the covariance of $\boldsymbol{G}\boldsymbol{\beta}$ with $\boldsymbol{G}\boldsymbol{\gamma}_{1}$ and $\boldsymbol{G}\boldsymbol{\gamma}_{2}$; $\rho_{1}$ and $\rho_{2}$ are the phenotypic correlations between $X$ and $Y_{1}$ and $Y_{2}$, respectively, in the population. The heritabilities $h^{2}_{y_{1},PB}$ and $h^{2}_{y_{2},PB}$ in the denominator are derived in (3). We note that $(1-\xi(\alpha)\rho^{2}_{1})$ and $(1-\xi(\alpha)\rho^{2}_{2})$ in the denominator can be regarded as the shrunk phenotypic variance of $Y_{1}$ and $Y_{2}$ after selection. The term $-\xi(\alpha)(\rho_{1}\rho_{G_{2}}+\rho_{2}\rho_{G_{1}})$ in the numerator is the first-order effects due to changes in the effect sizes of genetic variants from the population to the sample of participants. Furthermore, the term $\xi(\alpha)\rho_{G_{1}}\rho_{G_{2}}/(1-\xi(\alpha)h^{2}_{x})$ captures the effects due to the changes in the LD structure. The remaining terms in the numerator describe their combined effects.

### v3

@@ Abstract
It is increasingly recognized that participation bias can pose problems for genetic studies. Recently, to overcome the challenge that genetic information of non-participants is unavailable, it is shown that by comparing the IBD (identity by descent) shared and not-shared segments between participating relative pairs, one can estimate the genetic component underlying participation. That, however, does not directly address how to adjust estimates of heritability and genetic correlation for phenotypes correlated with participation. Here, for phenotypes whose mean differences between population and sample are known, we demonstrate a way to do so by adopting a statistical framework that separates the genetic and non-genetic correlations between participation and these phenotypes. Crucially, our method avoids making the assumption that the effect of the genetic component underlying participation is manifested entirely through these other phenotypes. Applying the method to 12 UK Biobank phenotypes, we found 8 that have significant genetic correlations with participation, including body mass index, educational attainment, and smoking status. For most of these phenotypes, without adjustments, estimates of heritability and the absolute value of genetic correlation would have underestimation biases.
@@ 1 Main
For sample surveys in general, a common approach to adjusting for ascertainment bias is to construct a propensity score based on variables, e.g. sex and educational attainment (EA), whose distributional differences between sample and target population are known, or can be well approximated 8. By assuming that the systematic component of participation probability is fully captured by this score, analyses can be adjusted by applying inverse propensity weighting (IPW) to the samples. For genetic studies, it has been shown that participation can be associated with many phenotypes, such as EA, alcohol use, mental and physical health 9; 10; 11; 12; 13; 14; 15. The UKBB participants are reported to be less likely to be obese, to smoke, to drink alcohol daily, and to have fewer self-reported health conditions compared with the general UK population 13. A recent study applied IPW to the UKBB data by creating a propensity score based on 14 phenotypes including age, BMI, weight, education, etc. 15 This propensity score, however, does not include genotypes. Thus, for analyses that involve genotypes, this adjustment would be sufficient only under the assumption that the systematic genotypic differences between sample and population are fully captured by this propensity score. This is equivalent to saying that the effect of the genetic component underlying participation is manifested entirely through this score, which can be considered as a composite phenotype. Previous results16 and results presented below show that this assumption does not hold.
@@
While the effectiveness of IPW or any other adjustment methods based only on phenotypes is doubtful, there is no obvious alternative unless genotype difference between sample and population can be independently estimated without relying on the phenotypes, which is difficult given that genotypes of non-participants are unavailable. A recent publication showed that the sample-population allelic frequency differences can be estimated by comparing the IBD (identity by descent) shared and not-shared segments among the participants 16. Here we utilize information obtained from that method through introducing a statistical model that specifies a genetic and a non-genetic component for each of the participation variable and each of the other phenotypes. The model separates the genetic and non-genetic correlations between participation and other correlated phenotypes, and allows us to obtain adjusted estimates of heritability and genetic correlation for these phenotypes that take PB into account. Biases of the estimated genetic components of the phenotypes and possible adjustments will also be discussed.
@@
Theoretically, without adjustments, heritability and genetic correlation can be over-estimated or under-estimated depending on the relative magnitudes of the genetic and non-genetic correlations between participation and the phenotypes. Empirically, we applied the adjustment method to 12 phenotypes of the UKBB data and found that, without adjustment, heritability and the absolute value of the genetic correlation estimates tend to be underestimated for most of them.
@@ 2.1 Participation model overview
In practice, participation is often a two-step process. In the first step, a group of individuals are invited to participate in the study, and in the second step the invited make the decision whether to participate. For simplicity, we consider a model where the bias is only in the second step, $i.e.$ the invited list is representative of the target population. Consequences of the violation of this assumption are discussed later. For the second step, as in Benonisdottir and Kong (2023)16, we adopt a liability-threshold model where the liability score of a person is denoted by $X$, which is assumed to have (approximately) a standard normal distribution in the population. An invited person participates if $X>t_{\alpha}$, where $t_{\alpha}=\Phi^{-1}(1-\alpha)$, $\Phi$ is the standard normal cumulative distribution function and $\alpha$ is the participation rate of those invited. A phenotype of interest is denoted by $Y$, standardized to have mean zero and variance 1. For an individual in the invited list, we assume it is a random draw following the following model:
$X=G_{x}+{\epsilon}_{x},$ (1)
$Y=G_{y}+{\epsilon}_{y},$
@@
where $G_{x}$ and $G_{y}$ are the genetic components of $X$ and $Y$, respectively. We assume that they can be represented by weighted sums of genotypes. Here ${\epsilon}_{x}$ and ${\epsilon}_{y}$ are parts of $X$ and $Y$ that are orthogonal (uncorrelated) to the genetic components $G_{x}$ and $G_{y}$, which could be partly random and partly determined by non-genetic factors. We denote the heritability of $X$ and $Y$ in the population as $h^{2}_{x}$ and $h^{2}_{y}$, which measure the phenotypic variance explained by genetic components. In model (1), we have $\mathrm{Var}(G_{x})=h^{2}_{x}$, $\mathrm{Var}(G_{y})=h^{2}_{y}$, and $\mathrm{Var}({\epsilon}_{x})=1-h^{2}_{x}$, $\mathrm{Var}({\epsilon}_{y})=1-h^{2}_{y}$. The genetic and non-genetic correlations between participation liability score $X$ and phenotype $Y$ in the population are denoted as $\rho_{g}=\mathrm{Corr}(G_{x},G_{y})$ and $\rho_{e}=\mathrm{Corr}({\epsilon}_{x},{\epsilon}_{y})$, respectively. The correlation between $X$ and $Y$, denoted by $\rho$, is equal to $\rho=\rho_{g}\sqrt{h^{2}_{x}h^{2}_{y}}+\rho_{e}\sqrt{(1-h^{2}_{x})(1-h^{2}_{y})}$.
To better illustrate the effect of PB, we reparametrize $G_{y}$ as $a\cdot G_{x}+G_{w}$, where $a=\rho_{g}\sqrt{h^{2}_{y}/h^{2}_{x}}$ and $G_{w}$ is orthogonal to $G_{x}$ with $\mathrm{Var}(G_{w})=(1-\rho^{2}_{g})h^{2}_{y}$. Similarly, ${\epsilon}_{y}$ is reformulated as $b\cdot{\epsilon}_{x}+{\epsilon}_{w}$, where $b=\rho_{e}\sqrt{(1-h^{2}_{y})/(1-h^{2}_{x})}$ and ${\epsilon}_{w}$ is orthogonal to ${\epsilon}_{x}$ with $\mathrm{Var}({\epsilon}_{w})=(1-\rho^{2}_{e})(1-h^{2}_{y})$. Figure 1 is a directed graph illustrating the mathematical causal relationships of the variables. The four explanatory variables $G_{x}$, $G_{w}$, ${\epsilon}_{x}$, and ${\epsilon}_{w}$ are independent of each other in the population. PB as a result of conditioning on $X>t_{\alpha}$, has the following impact on the sample. As $G_{w}$ and ${\epsilon}_{w}$ are not causal variables of $X$, their distributions are not affected by the conditioning, and they remain uncorrelated with each other and with $G_{x}$ and ${\epsilon}_{x}$. By contrast, the variances of $G_{x}$ and ${\epsilon}_{x}$ shrink, and furthermore they become negatively correlated because of collider bias.
Figure 1: Graphical representation of the contributions of genetic and non-genetic variables underlying the participation liability score $X$ and another phenotype $Y$. The genetic and non-genetic components of $Y$ are reformulated as $G_{y}=a\cdot G_{x}+G_{w}$ and ${\epsilon}_{y}=b\cdot{\epsilon}_{x}+{\epsilon}_{w}$. Variables $G_{x},G_{w},{\epsilon}_{x}$, and ${\epsilon}_{w}$ are orthogonal in the population. Each arrow is accompanied by the causal strength of the explanatory variable, quantified as the $r^{2}$ between the two variables. As $X$ and $Y$ are standardized, the contributions of the two arrows pointing to $X$ and the four arrows pointing to $Y$ each sum to 1. Upon conditioning on participation, $X>t_{\alpha}$, variances of $G_{x}$ and ${\epsilon}_{x}$ are attenuated and they become negatively correlated (the red dashed line) as a result of collider bias.
@@ 2.2 Heritability estimated with participation bias
Under model (1), the heritability of phenotype $Y$ can be calculated as $h^{2}_{y}=\mathrm{Corr}(Y,G_{x})^{2}$. Let
$h^{2}_{y,PB}=\mathrm{Corr}(Y,G^{\prime}_{y}\mid X>t_{\alpha})^{2},$ (2)
which is the “apparent” heritability that is estimated using the sample of participants without considering PB. Here $G^{\prime}_{y}$ is the genetic component in the sample of participants, and $h^{2}_{y,PB}$ can be regarded as the expected proportion of the phenotypic variance that can be explained by genetic variants alone in the sample. Specifically, $G^{\prime}_{y}$ can be written as $a^{\prime}\cdot G_{x}+G_{w}$ (expression for $a^{\prime}$ is given in the Supplementary Note 1.2). Consider here the more common case where $\rho_{g}$ and $\rho_{e}$ have the same sign. Suppose both $\rho_{g}$ and $\rho_{e}$ are positive, which implies $a$ and $b$ are positive (for cases where these parameters are all negative, the discussion here would apply to $-Y$). In this case, $a^{\prime}<a$ because $G_{x}$ and ${\epsilon}_{x}$ are negatively correlated upon conditioning, and thus the coefficient of $G_{x}$ is reduced because it is capturing some of the positive contribution of ${\epsilon}_{x}$ to $Y$ in a negative manner. Since $a^{\prime}\cdot G_{x}+G_{w}=G_{y}-(a-a^{\prime})\cdot G_{x}$, this means that the genetic component of $Y$ estimated from the sample is biased towards reducing the contribution of the $G_{x}$ component. After selection, the variance of $Y$ explained by $G_{w}$ remains unchanged, while, ignoring rare cases where $a^{\prime}<0$ and $|a^{\prime}|>a$, the variance accountable by $G_{x}$ would decrease as both its variance and its coefficient are reduced. The variance of $Y$ also shrinks because of reduced contributions from both $G_{x}$ and ${\epsilon}_{x}$. As both the variance accountable by genetic components and the total variance shrink, their ratio ($i.e.$ $h^{2}_{y,PB}$) can either increase or decrease depending on the relative shrinkages.
The quantitative results presented below are derived assuming that the genetic and non-genetic components ($G_{x},G_{w},{\epsilon}_{x},{\epsilon}_{w}$) jointly follow a multivariate normal distribution (MVN), which should be appropriate for complex traits. The expressions appear more complicated than usual with MVN manipulations because conditioning is on an interval, instead of on a single value, of $X$. Based on the properties of truncated MVN17; 18, we show that (see Supplementary Note 1.2)
$h^{2}_{y,PB}=\frac{1}{1-\xi(\alpha)\rho^{2}}\left[h^{2}_{y}-\xi(\alpha)\rho_{G}(\rho_{G}+2\rho_{E})+\frac{\xi(\alpha)^{2}h^{2}_{x}\rho_{E}^{2}}{1-\xi(\alpha)h^{2}_{x}}\right],$ (3)
where $\rho_{G}=\rho_{g}\sqrt{h^{2}_{x}h^{2}_{y}}$ is the covariance of $G_{x}$ and $G_{y}$; $\rho_{E}=\rho_{e}\sqrt{(1-h^{2}_{x})(1-h^{2}_{y})}$ is the covariance of ${\epsilon}_{x}$ and ${\epsilon}_{y}$; and
$\xi(\alpha)=\left[\frac{\phi(t_{\alpha})}{\alpha}\right]^{2}-t_{\alpha}\cdot\left[\frac{\phi(t_{\alpha})}{\alpha}\right],$ (4)
where $\phi$ is the density function of the standard normal distribution. Notably, $\left[1-\xi(\alpha)\right]=\mathrm{Var}(X|X>t_{\alpha})$ and the denominator of (3), $(1-\xi(\alpha)\rho^{2})$, is the variance of $Y$ after selection. The numerator of (3) is the variance of $Y$ that can be accounted for by $G_{x}$ and $G_{w}$ after selection.
@@
A comparison of $h^{2}_{y}$ and $h^{2}_{y,PB}$ is shown in Figure 2(a). In general, with other parameters fixed, the difference between $h^{2}_{y}$ and $h^{2}_{y,PB}$ increases as $\alpha$ decreases. As noted earlier, PB can lead to both upward and downward biases for the estimation of heritability in the sample of participants. For example, if the correlation between $X$ and $Y$ arises solely from genetic components, conditioning on participation will reduce the variance of the genetic component of $Y$ in the selected sample. With the variance of the non-genetic component unchanged, the proportion of the genetic effects in the sample will shrink. Conversely, if the correlation between the phenotype and participation is induced by non-genetic factors only, the proportion of variance of $Y$ attributed to the genetic component in the sample will increase, which leads to an upward bias in the estimation of heritability.
@@
Figure 2: Theoretical results for heritability and genetic correlation in the sample under various genetic architectures. (a) The heritability of a phenotype of interest other than participation. (b) The genetic correlation between two phenotypes (other than participation). (c) The genetic correlation between the liability score of participation and the phenotype. The heritability of participation liability score and other phenotypes are fixed at 0.125 and 0.5. Dashed lines indicate population value and blue solid lines indicate what is being estimated based on the biased sample without adjustment.
@@ 2.3 Genetic correlation estimated with participation bias
Here we consider two phenotypes, $Y_{1}$ and $Y_{2}$, both following the model (1), and $\varphi_{g}$ denotes their genetic correlation in the population. The corresponding genetic covariance is $\varphi_{G}=\varphi_{g}\sqrt{h^{2}_{y_{1}}h^{2}_{y_{2}}}$, where $h^{2}_{y_{1}}$ and $h^{2}_{y_{2}}$ are the heritability of the two phenotypes. Let $\varphi_{g,PB}=\mathrm{Corr}(G^{\prime}_{y_{1}},G^{\prime}_{y_{2}}|X>t_{\alpha})$. In parallel with $h^{2}_{y,PB}$, the quantity $\varphi_{g,PB}$ represents the “apparent” correlation between the genetic components of $Y_{1}$ and $Y_{2}$ that is being estimated ignoring PB. We show in Supplementary Note 1.3 that,
$\varphi_{g,PB}=\frac{\varphi_{G}-\xi(\alpha)(\rho_{E_{1}}\rho_{G_{2}}+\rho_{E_{2}}\rho_{G_{1}}+\rho_{G_{1}}\rho_{G_{2}})+\frac{\xi(\alpha)^{2}h^{2}_{x}\rho_{E_{1}}\rho_{E_{2}}}{1-\xi(\alpha)h^{2}_{x}}}{\sqrt{(1-\xi(\alpha)\rho^{2}_{1})(1-\xi(\alpha)\rho^{2}_{2})}\cdot\sqrt{h^{2}_{y_{1},PB}h^{2}_{y_{2},PB}}}.$ (5)
Here $\rho_{G_{1}}$ and $\rho_{G_{2}}$ are the covariance of $G_{x}$ with $G_{y_{1}}$ and $G_{y_{2}}$; $\rho_{E_{1}}$ and $\rho_{E_{2}}$ are the covariance of ${\epsilon}_{x}$ with ${\epsilon}_{y_{1}}$ and ${\epsilon}_{y_{2}}$; $\rho_{1}$ and $\rho_{2}$ are the phenotypic correlations between $X$ and $Y_{1}$ and $Y_{2}$, respectively, in the population. The denominator is the squared root of the product of variances explained by $G^{\prime}_{y_{1}}$ and $G^{\prime}_{y_{2}}$ in the sample of participants, where the heritabilities $h^{2}_{y_{1},PB}$ and $h^{2}_{y_{2},PB}$ are derived in (3). The numerator is the covariance between $G^{\prime}_{y_{1}}$ and $G^{\prime}_{y_{2}}$, which incorporates the negative correlations between $G_{x_{1}}$ and ${\epsilon}_{x_{1}}$, and $G_{x_{2}}$ and ${\epsilon}_{x_{2}}$ in the sample of participants.